OpenStudy (kuoministers):

show that each expression is rational question will be posted below HELP MEDALS WILL BE GIVEN

OpenStudy (kuoministers):

\[\frac{ 3 }{ 2 +\sqrt{2} } + \frac{ 3 }{ \sqrt{2} }\] what does rational even mean :P

OpenStudy (anonymous):

Rational means that it can be written as a fraction where the numerator and denominator are both integers.

OpenStudy (kuoministers):

mhm ok... but how can u prove it is rational or show it

OpenStudy (anonymous):

multiply the top and bottom of the first term by 2-sqrt(2) and the top and bottom of the second term by sqrt(2). What happens?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2+\sqrt{2} }+\frac{ 3 }{ \sqrt{2} }\] \[=\frac{ 3\sqrt{2} + 3(2+\sqrt{2}) }{ \sqrt{2}(2+\sqrt{2}) }\] \[=\frac{ 6\sqrt{2}+6 }{ 2\sqrt{2}+2 }\] \[=(\frac{ 6\sqrt{2}+6 }{ 2\sqrt{2}+2 })\times(\frac{ 2\sqrt{2} - 2 }{ 2\sqrt{2} - 2 })\] \[=\frac{ (6\sqrt{2}+6)(2\sqrt{2} - 2) }{ (2\sqrt{2}+2)(2\sqrt{2} - 2) }\] Expand the numerator as per usual, and then use your knowledge of completing the square in the denominator. \[=\frac{ 12(2)-12\sqrt{2} + 12\sqrt{2} -12 }{ 4(2) - 4 }\]

OpenStudy (anonymous):

All the terms with square root signs cancel out. Therefore you have a rationalised term. \[=\frac{ 12 }{ 4 }\]

OpenStudy (kuoministers):

wow thanks so much